Business and finance. The demand equation for a certain type of printer is given by D=-200p + 35,000

The supply equation is predicted to be s=-p^+400p-20,000

Find the equilibrium price.

Answer by stanbon(29603)

The demand equation for a certain type of printer is given by D=-200p + 35,000
The supply equation is predicted to be s=-p^+400p-20,000
-------------------------
D = S
-200p+35000 = -p^2+400p-20000
p^2-600p+55000 = 0
p = [600 +- sqrt(600^2-4*55000)]/2
p = [600 +- 374.1657]/2
p = 974.1657/2 or x = 225.8343/2
p = 487.08 or p = 112.92
===========================
Cheers,
Stan H.